OpenStudy (anonymous):

121+x^2=225 what does x equal?

OpenStudy (anonymous):

so you want to get the x on its own. try subtracting 121 from both sides first

OpenStudy (anonymous):

104

OpenStudy (anonymous):

yeah, x^2 = 104

OpenStudy (anonymous):

now you need to get rid of the ^2

OpenStudy (anonymous):

mmk how do i do that?

OpenStudy (anonymous):

what is the opposite of ^2?

OpenStudy (anonymous):

okay well whats the opposite of squaring something?

OpenStudy (anonymous):

remember it's \[\pm \sqrt{104}\]

OpenStudy (anonymous):

ooh nevermind went blank i would square root it making x = 10.2 right

OpenStudy (anonymous):

121+x^2=225 -121 -121 x^2=104 square root x square root 104 X=10.2 approximately

OpenStudy (anonymous):

yeah awesome :)

OpenStudy (anonymous):

=) thank-you guys!!

OpenStudy (anonymous):

but it's gonna be a positive or negative 10.2

OpenStudy (anonymous):

Aww man! she got it! oh well, congratulations! That's how you would show your work.

OpenStudy (anonymous):

positive right...

OpenStudy (anonymous):

lol thanks =)

OpenStudy (anonymous):

It can actually be positive and negative!

OpenStudy (anonymous):

could be either, but i doubt your teacher will care at this stage

OpenStudy (anonymous):

it can be positive and negative check it wit the calculator yu shud use positive for now

OpenStudy (anonymous):

haha yeahh i got it right =) thanks again all of you

OpenStudy (anonymous):

You are welcome!